OpenStudy (anonymous):

Write a quadratic function in standard form having zeros of -5 and 1

OpenStudy (anonymous):

zeros means its roots are -5 and 1 Forming quad eq x^2-(sum of roots)x +(product of roots) = 0

OpenStudy (anonymous):

So h(x)=x^2+4x-4

OpenStudy (anonymous):

Product = -5*1 = -5

OpenStudy (anonymous):

if u r satisfied give medals by clicking best answer option